# Counterfactual Debating with Preset Stances for Hallucination Elimination of LLMs

### v1

@@ Abstract
Yi Fang1, Moxin Li2, Wenjie Wang2thanks: Corresponding author., Hui Lin3, Fuli Feng [^1] footnotemark: 1 1University of Science and Technology of China, 2National University of Singapore, 3Electronic Science Research Institute of China Electronics peterfang@mail.ustc.edu.cn, limoxin@u.nus.edu, {wenjiewang96,fulifeng93}@gamil.com, linhui@whu.edu.cn
@@ 1 Introduction
To this end, we propose a CounterFactual Multi-Agent Debate (CFMAD) framework comprising two key stages: abduction generation and counterfactual debate. In the abduction generation stage, LLMs are tasked with producing potential correct reasons for a predetermined answer. Subsequently, in the counterfactual debate stage, a structured debate method is employed to assess these abductions and ascertain the sole correct response. Specifically, we introduce a critic who questions the validity of each generated abduction, and prompt the LLM to defend its position in a debate with the critic. The deliberation is then presented to an impartial third-party judge for final adjudication. Extensive experiments spanning fact-checking, reading comprehension, and commonsense reasoning tasks validate the effectiveness of CFMAD over existing benchmarks across four datasets. We release our code and data at https://anonymous.4open.science/r/CFMAD-468D/.
@@ 3 Counterfactual Multi-agent Debate
To address the overconfidence issue, the key lies in overriding the inherent biases and beliefs of LLMs for answer generation. To achieve this, we consider initially configuring the LLMs with various stances, allowing them to hypothesize the correctness of each possible answer and uncover the underlying rationale of each answer. This approach compels LLMs to inspect all potential answers, liberating them from inherent biases and beliefs. Subsequently, we can critically assess the potential rationales to identify the correct answer.
@@ 3.2 Counterfactual Debate
After multi-round debating, we present the debate process of all agent-critic pairs to a third-party judge, enabling them to meticulously analyze and juxtapose the varied debate trajectories, thereby discerning the final answer.
@@ 4.1 Experimental Setup
We conduct experiments on four datasets: Hover (Jiang et al., 2020), BoolQ (Clark et al., 2019), CosmosQA (Huang et al., 2019), and CommonsenseQA (Talmor et al., 2019). Hover and BoolQ are binary prediction tasks with only true or false answers. CosmosQA and CommenseQA are multi-choice tasks with 4 and 5 options, respectively. Note that we split Hover into two subsets named Hover 3-hop and Hover 4-hop with questions requiring 3 and 4 steps of reasoning, respectively. Comparison between Hover 3-hop and Hover 4-hop might reveal the influence of problem difficulty on method effectiveness since more complex questions typically require more reasoning hops. More details about these datasets can be found in the Appendix A.1.
@@
The implementation detail involves three key factors: backbone, prompt, and inference temperature. For all compared methods, we use GPT-3.5-turbo-0613111https://chatgpt.com/. as our backbone LLM and present their prompts in Appendix B and C. As to the inference temperature, we set it to 0.2 in most methods for the sake of fair comparison. The only exception is Self-Consistency, where we follow the original paper and set the temperature to 1 since the method requires high diversity of samples (Wang et al., 2023).
@@ 6 Conclusion
Appendix A Experiments Details
A.1 Dataset Details
@@
A.2 Method Implementation Details
@@
Appendix B Basline Prompts
B.1 CoT Prompt
@@
    Output format:
@@ B.2 Relfection Prompt
    Assistant’s reply: {CoT_reply}
@@
    Assistant’s reply: {CoT_reply}
    Feeback: {reflection_reply}
@@ B.3 MAD
    Assistant: {Your previous response}
    Other agent1: {Other agents’ previous responses1}
    Other agent2: {Other agents’ previous responses2}
@@
    Agent1: {last response of agent 1}
    Agent2: {last response of agent 2}
    Agent3: {last response of agent 3}
@@ 4. Format Requirement: Each prompt instruction should start with ### and end with @@@
    The input question is: {question}. Please generate the most suitable three prompts:
@@ 3. Why are the answers of the two judgements different?
    Judgement1: {reply1},
    Judgement2: {reply2},
    Judgement3: {reply3}
@@
    “Judgement”: “{reply1}”,
    “Judgement”: “{reply2}”,
    “Judgement3”: “{reply3}”
@@
    “relation”: {difference_2_3}
    Checklist: {checklist}
@@
Appendix C Our Prompts
C.1 Fact Check Task
@@
    Claim: {claim}
    Please fully understand the evidence and claim, and answer why the claim is {true/false}?
@@
    Claim: {claim}
    Assistant: {reply of assistant}
@@
    Fact checker: {reply of assistant}
    Critic: {reply of crtic}
@@ C.2 Commonsense Reasoning
    Try to explain why the question’s answer might be option {predetermined answer}.
@@
    Judgement: The answer is option {predetermined answer}.
@@
    Assistant: {reply of assistant}
@@
    Assistant: {reply of assistant}
    Critic: {reply of critic}

### v2

@@ Abstract
Yi Fang1, Moxin Li2, Wenjie Wang2thanks: Corresponding author., Hui Lin3, Fuli Feng [^1] footnotemark: 1 1University of Science and Technology of China, 2National University of Singapore, 3Electronic Science Research Institute of China Electronics peterfang@mail.ustc.edu.cn, limoxin@u.nus.edu, {wenjiewang96,fulifeng93}@gmail.com, linhui@whu.edu.cn
@@ 1 Introduction
To this end, we propose a CounterFactual Multi-Agent Debate (CFMAD) framework comprising two key stages: abduction generation and counterfactual debate. In the abduction generation stage, LLMs are tasked with producing potential correct reasons for a predetermined answer. Subsequently, in the counterfactual debate stage, a structured debate method is employed to assess these abductions and ascertain the sole correct response. Specifically, we introduce a critic who questions the validity of each generated abduction, and prompt the LLM to defend its position in a debate with the critic. The deliberation is then presented to an impartial third-party judge for final adjudication. Extensive experiments spanning fact-checking, reading comprehension, and commonsense reasoning tasks validate the effectiveness of CFMAD over existing benchmarks across four datasets. We release our code and data at https://github.com/Peter-Fy/CFMAD/.
@@ 3 Counterfactual Multi-agent Debate
To address the overconfidence issue, the key lies in overriding the inherent biases and beliefs of LLMs for answer generation. To achieve this, we consider initially configuring the LLMs with various stances, allowing them to hypothesize the correctness of each possible answer and uncover the underlying rationale of each answer. This approach compels LLMs to inspect all potential answers, reducing their reliance on inherent biases and beliefs. Subsequently, we can critically assess the potential rationales to identify the correct answer.
@@ 3.2 Counterfactual Debate
After multi-round debating, we present the debate process of all agent-critic pairs to a third-party judge, implemented as another LLM. This enables the judge to meticulously analyze and juxtapose the varied debate trajectories, thereby discerning the final answer. An example prompt for the third-party judge can be found in the Appendix D.
@@ 4.1 Experimental Setup
We conduct experiments on four datasets: Hover (Jiang et al., 2020), BoolQ (Clark et al., 2019), CosmosQA (Huang et al., 2019), and CommonsenseQA (Talmor et al., 2019). Hover and BoolQ are binary prediction tasks with only true or false answers. CosmosQA and CommenseQA are multi-choice tasks with 4 and 5 options, respectively. Note that we split Hover into two subsets named Hover 3-hop and Hover 4-hop with questions requiring 3 and 4 steps of reasoning, respectively. Comparison between Hover 3-hop and Hover 4-hop might reveal the influence of problem difficulty on method effectiveness since more complex questions typically require more reasoning hops. More details about these datasets can be found in the Appendix B.1.
@@
The implementation detail involves three key factors: backbone, prompt, and inference temperature. For all compared methods, we use GPT-3.5-turbo-0613111https://chatgpt.com/. as our backbone LLM and present their prompts in Appendix C and D. As to the inference temperature, we set it to 0.2 in most methods for the sake of fair comparison. The only exception is Self-Consistency, where we follow the original paper and set the temperature to 1 since the method requires high diversity of samples (Wang et al., 2023).
@@ 6 Conclusion
Acknowledgments
We thank the reviewers for their constructive feedback.
Appendix A Complementary Experiments
A.1 Token Efficiency Analysis
We compared the token costs of our method to several baselines on the CommonsenseQA dataset, as shown in Table 3. Specifically, We computed the number of prompts used per data sample, the average number of tokens per prompt for a single sample, and the total average token usage per sample. The results indicate that our method’s token costs are similar to those of other multi-agent methods, such as MAD and Self-Contrast. Similar results are observed on other datasets.
Method Prompts/Sample Tokens/Prompt Tokens/Sample
CoT 1 103.25 103.25
Self-Reflection 3 255.86 767.59
Self-Consistency 7 103.25 722.73
MAD 10 276.72 2767.18
Self-Contrast 6 349.80 2098.81
CFMAD (Ours) 10 246.82 2468.22
Table 3: Token Costs of different methods.
A.2 Results on Open-Ended Tasks
In addition to evaluating binary prediction and multi-choice QA tasks, where candidate answers are provided, we also conducted evaluations on more open-ended tasks, including mathematical reasoning and machine translation. Specifically, we selected three mathematical reasoning datasets—SVAMP (Patel et al., 2021), GSM8K (Cobbe et al., 2021), and Multi_Arith (Roy and Roth, 2016)—and one machine translation dataset, CommonMT (He et al., 2020). These datasets do not offer predefined answer choices and require the language models (LLMs) to generate open-ended responses.
Since these tasks lack candidate answers, we first used CoT prompts to generate candidate answers from the open answer spaces by sampling multiple times. This allowed us to transform the problems into a multi-choice QA format, which could then be processed using our pipeline.
For evaluation, we used precision as a metric for the mathematical reasoning tasks and BLEURT1 for the machine translation task. The results, as shown in the Table 4, demonstrate that CFMAD outperforms the other baselines, highlighting the effectiveness and generality of our approach.
Method SVAMP GSM8K Multi_Arith CommonMT
CoT 0.8067 0.7567 0.9575 0.7008
Self-Consistency 0.8500 0.8167 0.9804 /
MAD 0.8567 0.8100 0.9575 0.7183
Self-Contrast 0.8467 0.8133 0.9771 0.7119
CFMAD 0.8700 0.8300 0.9837 0.7201
Table 4: Results on Open-Ended Tasks
A.3 Results on Open-Sourced LLMs
In addition to GPT-3.5, we have also assessed our method on two other open-source LLMs: Llama2-7B (Touvron et al., 2023) and Llama3-8B (Dubey et al., 2024). The results on the CommonsenseQA dataset are presented in Table 5. Our approach, CFMAD, consistently outperforms other baselines across these models, highlighting the effectiveness of our method when applied to open-source LLMs.
CommonsenseQA Llama2-7B Llama3-8B
Self-Consistency 0.6233 0.7533
MAD 0.5900 0.7467
Self-Contrast 0.6233 0.7600
CFMAD 0.6367 0.7700
Table 5: Results on Open-Sourced Tasks
Appendix B Experiments Details
B.1 Dataset Details
@@
B.2 Method Implementation Details
@@
Appendix C Basline Prompts
C.1 CoT Prompt
@@
    Play the role of a common sense reasoning expert. Choose the most appropriate answer for the question. You are expected to explain your reasoning process step-by-step before providing the final answer. Output format:
@@ C.2 Relfection Prompt
    Assistant’s reply:
    {CoT_reply}
@@
    Assistant’s reply:
    {CoT_reply}
    Feeback:
    {reflection_reply}
@@ C.3 MAD
    Assistant:
    {Your previous response}
    Other agent1:
    {Other agents’ previous responses1}
    Other agent2:
    {Other agents’ previous responses2}
@@
    Agent1:
    {last response of agent 1}
    Agent2:
    {last response of agent 2}
    Agent3:
    {last response of agent 3}
@@ 4. Format Requirement: Each prompt instruction should start with ### and end with @@@
    The input question is:
    {question}. Please generate the most suitable three prompts:
@@ 3. Why are the answers of the two judgements different?
    Judgement1:
    {reply1},
    Judgement2:
    {reply2},
    Judgement3:
    {reply3}
@@
    “Judgement”: “
    {reply1}”,
    “Judgement”: “
    {reply2}”,
    “Judgement3”: “
    {reply3}
@@
    “relation”:
    {difference_2_3}
    Checklist:
    {checklist}
@@
Appendix D Our Prompts
D.1 Fact Check Task
@@
    Claim:
    {claim}
    Please fully understand the evidence and claim, and answer why the claim is
    {true/false}?
@@
    Claim:
    {claim}
    Assistant:
    {reply of assistant}
@@
    Fact checker:
    {reply of assistant}
    Critic:
    {reply of crtic}
@@ D.2 Commonsense Reasoning
    Try to explain why the question’s answer might be option
    {predetermined answer}.
@@
    Judgement: The answer is option
    {predetermined answer}.
@@
    Assistant:
    {reply of assistant}
@@
    Assistant:
    {reply of assistant}
    Critic:
    {reply of critic}